Replacement times for TV sets has a mean of 8.2 years and a standard deviation of 1.1 years.


a. What is the probability that a randomly selected TV set will have replacement time in less than 8 years? Give your answer to four decimal places.
Answer = Answer


b. What is the probability that 40 randomly selected TV sets will have a mean replacement time that is less than 8 years? Give your answer to four decimal places.
Answer = Answer

Answer by ewatrrr(24785)


Hi
Normal distribution:    μ = 8.2years   σ = 1.1years 
Using TI or similarly an inexpensive calculator like a Casio fx-115 ES plus

a. P(x < 8) = normalcdf(-9999, 8,8.2,1.1) = 0.4279

b. P( x̄ < 8)= normalcdf(-9999, 8,8.2,1.1/√40) = .125
or
z = (8-8.2)/1.1/sqrt(40) = -1.1499 
P( z ≤ -1.1499) = .125

Wish You the Best in your Studies.